OpenStudy (anonymous):

locate the critical points and classify them as stationary points or points of nondifferentiability for f(x)=cos3x

OpenStudy (anonymous):

derivative is \(-3\sin(3x)\) zeros of the derivative are where \(3x=0, \pi, 2\pi,...\)

OpenStudy (anonymous):

-so \(x=\frac{k\pi}{3}\)

OpenStudy (anonymous):

they are clearly stationary points, since cosine is infinitely differentiable

OpenStudy (anonymous):

oh okay i see thank you :)